OpenStudy (rissa):

Find the amount of decrease and the percent decrease if the original amount is 65 and the new amount is -780.

OpenStudy (rissa):

A. 845; 1310% B. 845; 1300% C. 845; 1295% d. 845; 1400%

OpenStudy (anonymous):

how much is the decrese????

OpenStudy (anonymous):

from 65 to -780 ?????

OpenStudy (rissa):

thats all the info it gives me :\

OpenStudy (anonymous):

65-845 = -780

OpenStudy (anonymous):

so the original number is decresed by 845

OpenStudy (anonymous):

now 845 is what percent of 65 ??????

OpenStudy (anonymous):

u dont know to calculate percent ??????

OpenStudy (anonymous):

\[\frac{845 \times 100}{65} = 1300\]